August 24, 1999 Day Shift

Taku Yamanaka (taku@fnal.gov)
Tue, 24 Aug 1999 15:30:08 -0500

Shift Summary for Day Shift, Aug. 24, 1999

On Shift: Erik Ramberg, Taku Yamanaka

Hours of beam available at all intensities: 2h
Hours of beam available at nominal intensity: 2h
Hours of E832 data (special runs not included): 1.3h
Hours of special runs (laser scans, muon runs...): 0.7

Accounting of lost time:

20min : KFREND hung up, fresh start
20min : Beam stop closed.

========================================================================
Runs ended this shift:

1310, 14311, 14312, 14313

=======================================================================

Tapes taken to FCC this shift

========================================================================
Accesses during shift:

Sasha G made opened and closed TRD gas valves.

========================================================================
Overview of shift:

Beam down from 9:00 - 3:30.
Beam intensity flucutated during spill.
Beam stop kicked in several times.

========================================================================
Description of Shift:

7:00 Beginning of day shift.

7:10 sp 256 BS kicked in, called MCR.
sp 260 BS opened.

7:27 sp 268 BS kicked in.
7:29 Vertical beam position is off. Called MCR.
7:44 sp 281 BS is out.

8:05 ktrig panel died, when trying to do KFREND.
Initialized frontend from text version of kfrend.
-- BAD idea. Since the window is still bad,
zapped ktrig, and restared ktrig.tcl.
8:18 sp 11 Yes, noticed that spill number got reset.
Stop Run 14310, fresh start.

8:33 Start Run 14311.
8:52 sp 16 Spill structure (DC1 current) is really bad. Fluctuating
by more than twice. Called MCR.

9:01 sp 23 Since there will be accelerator studies, BS is closed.
9:03 Stopped Run 14311.

9:08 MCR called. They will test BS.

9:17 Start Run 14312, LASER_SCAN.

~10AM Run 14312 ended.

10:14 Start Run 14313 Cosmics.
10:50 Stop Run 14313.

14:50 Sasha G made an access with Taku, and closed TRD gas
valves.

========================================================================
Follow up Notes on Previous Shift

Watch out for the beam structure (DC1 current scope). They are supposed
to have worked on it (QXR), so it should be better.

Watch the beam stop. It had been kicking in often this morning.

Watch the beam position.